what is x-intercept of the line passing through points (-4,15) and having a slope of -1?
Answer by lwsshak3(11628)

what is x-intercept of the line passing through points (-4,15) and having a slope of -1
equation of a line in slope-intercept form: y=mx+b, m=slope,b=y-intercept
y=-x+b
solve for b using given coordinates(4,15) on the line
b=y-x=15-4=11
equation: y=-x+11
y-intercept=11